# Tuning-Free Structured Sparse PCA via Deep Unfolding Networks

### v1

@@ Abstract
00footnotetext: This work was supported by the National Natural Science Foundation of China under Grant 12371306 (Corresponding author: Xianchao Xiu.)
@@ 1 Introduction
Unsupervised feature selection (UFS) has emerged as a critical technique in high-dimensional data analysis, particularly for image and signal processing applications where class labels are often unavailable; see [1] for a recent survey. Among various approaches, sparse principal component analysis (PCA) [2] has become a prominent UFS tool, demonstrating remarkable success across diverse domains ranging from image processing [3], fault diagnosis [4] to natural language processing [5].
@@
where $X\in\mathbb{R}^{d\times m}$ is the projection matrix and $I\in\mathbb{R}^{m\times m}$ is the identity matrix. However, it suffers from limited feature interpretability. To address this issue, sparse PCA was proposed in [6], by defining
@@
On the one hand, the $\ell_{1}$-norm in sparse PCA only exploits element-wise sparsity, which often lead to biased solutions. This limitation becomes particularly apparent in applications where features naturally exhibit group-wise sparsity, which cannot be characterized by $\ell_{1}$-norm but could be captured by $\ell_{2,1}$-norm (defined as the sum of $\ell_{2}$-norm of each row) [7]. Current methodologies usually consider these sparsity-inducing norms separately or apply them to different variables, while it remains an open question whether the integration of $\ell_{1}$-norm and $\ell_{2,1}$-norm with PCA could yield superior performance in the field of UFS.
On the other hand, traditional parameter selection strategies (e.g., grid search, Bayesian optimization) face significant scalability limitations in high-dimensional settings. The emerging paradigm of deep unfolding networks [8] offers a promising alternative by embedding optimization mechanics into neural architectures through algorithmic unfolding, i.e., transforming the iterative algorithm steps into network layers. This approach combines mathematical interpretability with neural network efficiency, achieving excellent performance in many fields such as communications [9], compressive sensing [10], and infrared small target detection [11]. Notably, deep unfolding networks typically require fewer trainable parameters than conventional deep architectures, which substantially reduces the computational complexity and training costs [12]. Despite these advancements, the application of deep unfolding networks to PCA-based UFS remains absent from the literature.
@@ 2 Methodology
This section describes how to leverage deep unfolding networks to develop a tuning-free algorithm for problem (3). Technically, we unfold each step of the alternating direction method of multipliers (ADMM) into a layer in neural networks, thus transforming parameters into learnable ones.
@@ 2.2 Update $Y^{k+1}$
$Y^{k+1}=\textbf{GSoftNet}_{\lambda/\alpha}(X^{k+1}+\Lambda^{k}/\alpha).$ (14)
@@
$Y^{k+1}=\textbf{GSoftNet}_{\lambda/\alpha}(X^{k+1}+\Lambda^{k}/\alpha)$
@@
$Z^{k+1}=\textbf{SoftNet}_{\mu/\beta}(X^{k+1}+\Pi^{k}/\beta)$
@@ 2.3 Update $Z^{k+1}$
$\frac{X^{k+1}+\Pi^{k}/\beta}{|X^{k+1}+\Pi^{k}/\beta|}\text{ReLU}(|X^{k+1}+\Pi^{k}/\alpha|-\mu/\alpha),$ (16)
@@
$Z^{k+1}=\textbf{SoftNet}_{\mu/\beta}(X^{k+1}+\Pi^{k}/\beta).$ (17)
@@ 4 Conclusion
In the future, we are interested in extending it to deep equilibrium networks to further improve the performance.

### v2

@@ Abstract
00footnotetext: This work was supported by the National Natural Science Foundation of China under Grant 12371306. (Corresponding author: Xianchao Xiu.)
@@ 1 Introduction
Unsupervised feature selection (UFS) has emerged as a critical technique in high-dimensional data analysis, particularly for image and signal processing applications where class labels are often unavailable; see [1, 2, 3] for a recent survey. Among various approaches, sparse principal component analysis (PCA) [4] has become a prominent UFS tool, demonstrating remarkable success across diverse domains ranging from image processing [5, 6], fault diagnosis [7, 8] to natural language processing [9, 10].
@@
where $X\in\mathbb{R}^{d\times m}$ is the projection matrix and $I\in\mathbb{R}^{m\times m}$ is the identity matrix. However, it suffers from limited feature interpretability. To address this issue, sparse PCA was proposed in [11] by defining
@@
On the one hand, the $\ell_{1}$-norm in sparse PCA only exploits element-wise sparsity, which often lead to biased solutions [12]. This limitation becomes particularly apparent in applications where features naturally exhibit group-wise sparsity, which cannot be characterized by $\ell_{1}$-norm but could be captured by $\ell_{2,1}$-norm (defined as the sum of $\ell_{2}$-norm of each row) [13, 14]. Current methodologies usually consider these sparsity-inducing norms separately or apply them to different variables, while it remains an open question whether the integration of $\ell_{1}$-norm and $\ell_{2,1}$-norm with PCA could yield superior performance in the field of UFS.
On the other hand, traditional parameter selection strategies (e.g., grid search, Bayesian optimization) face significant scalability limitations in high-dimensional settings [15, 16, 17]. The emerging paradigm of deep unfolding networks [18] offers a promising alternative by embedding optimization mechanics into neural architectures through algorithmic unfolding, i.e., transforming the iterative algorithm steps into network layers [19, 20, 21]. This approach combines mathematical interpretability with neural network efficiency, achieving excellent performance in many fields such as communications [22], compressive sensing [23], and infrared small target detection [24]. Notably, deep unfolding networks typically require fewer trainable parameters than conventional deep architectures, which substantially reduces the computational complexity and training costs [25, 26]. Despite these advancements, the application of deep unfolding networks to PCA-based UFS remains absent from the literature.
@@ 2 Methodology
This section describes how to leverage deep unfolding networks to develop a tuning-free algorithm for problem (3). Technically, we unfold each step of the alternating direction method of multipliers (ADMM) [28] into a layer in neural networks, thus transforming parameters into learnable ones.
@@ 2.2 Update $Y^{k+1}$
$Y^{k+1}=\textbf{{GSoftNet}}(X^{k+1}+\Lambda^{k}/\alpha,\lambda/\alpha).$ (14)
@@
$Y^{k+1}=\textbf{{GSoftNet}}(X^{k+1}+\Lambda^{k}/\alpha,\lambda/\alpha)$
@@
$Z^{k+1}=\textbf{{SoftNet}}(X^{k+1}+\Pi^{k}/\beta,\mu/\beta)$
@@ 2.3 Update $Z^{k+1}$
$\frac{X^{k+1}+\Pi^{k}/\beta}{|X^{k+1}+\Pi^{k}/\beta|}\text{ReLU}(|X^{k+1}+\Pi^{k}/\beta|-\mu/\beta),$ (16)
@@
$Z^{k+1}=\textbf{{SoftNet}}(X^{k+1}+\Pi^{k}/\beta,\mu/\beta).$ (17)
@@ 3.2 Numerical Results
3.3 Ablation Studies
Table 4 lists the clustering results without and with the network, where $\times$ means that the regularization parameters are selected by grid search. It can be seen that deep unfolding networks can significantly improve ACC and NMI. In fact, In fact, it is quite difficult for grid search to find out the proper regularization parameters. In addition, Table 5 lists the clustering results without and with dynamic parameters, where $\times$ indicates that the regularization parameters are fixed during the network iterations. Obviously, the dynamic parameters can better adapt to the data. The above ablation experiments validate the necessity of our method.
Table 4: Ablation studies for the network.
Datasets Network ACC NMI
COIL20 $\times$ 55.12$\pm$2.67 70.44$\pm$1.37
$\surd$ 57.46$\pm$2.76 72.21$\pm$2.68
Isolet $\times$ 51.84$\pm$2.82 67.02$\pm$1.43
$\surd$ 58.43$\pm$4.31 71.80$\pm$1.59
UMIST $\times$ 40.65$\pm$2.29 55.88$\pm$1.62
$\surd$ 47.58$\pm$4.97 66.62$\pm$7.52
MSTAR $\times$ 80.65$\pm$6.47 80.53$\pm$2.41
$\surd$ 81.90$\pm$6.87 80.67$\pm$3.47
Table 5: Ablation studies for dynamic parameters.
Datasets Dynamic ACC NMI
COIL20 $\times$ 56.71$\pm$3.83 71.49$\pm$3.67
$\surd$ 57.46$\pm$2.76 72.21$\pm$2.68
Isolet $\times$ 52.06$\pm$3.71 68.91$\pm$2.36
$\surd$ 58.43$\pm$4.31 71.80$\pm$1.59
UMIST $\times$ 42.63$\pm$2.78 60.12$\pm$1.69
$\surd$ 47.58$\pm$4.97 66.62$\pm$7.52
MSTAR $\times$ 80.74$\pm$5.28 80.59$\pm$3.67
$\surd$ 81.90$\pm$6.87 80.67$\pm$3.47
Figure 4: Visualization of deep unfolding stages on COIL20.
3.4 Remark
Compared with traditional numerical optimization, which often requires hundreds or thousands of iterations, deep unfolding networks only require less than 10 stages to achieve good results; see Fig. 4. However, it should be noted that the performance will degrade if the number of iterations exceeds a certain number, so in practice we chose between 3 and 5.
@@ 4 Conclusion
In the future we are interested in designing more effective loss functions and extending it to deep equilibrium networks [36] to further improve the performance.